Assume 5ft/s of Hydrogen at 15 psia and 80F are compressed during an irreversible adiabatic process to 90 psia and 522F. For a non flow process determine a)k b)V c) W d)U and e)s


1
Expert's answer
2021-04-01T01:10:12-0400

"T_2 = 522°F = 545.5K
\\ T_1 = 80°F= 300K"


"P_1= 15psia\\
P_2 =90psia"


∆T = 522F -80F = 442F = 228K

∆P = 90psia - 15psia = 75psia


speed rate = 5ft/s


"\dfrac{V_1}{V_2} = \dfrac{P_2×T_1}{P_1×T_2} = \dfrac{90 × 300}{15× 545.5} = 3.3"


Therefore the final volume is 3.3 times lesser than the initial volume.